# Predict. Optimize. Revise. On Forecast and Policy Stability in Energy Management Systems

### v4

@@ Abstract
This research addresses the challenge of integrating forecasting and optimization in energy management systems, focusing on the impacts of switching costs, forecast accuracy, and stability. It proposes a novel framework for analyzing online optimization problems with switching costs and enabled by deterministic and probabilistic forecasts. Through empirical evaluation and theoretical analysis, the research reveals the balance between forecast accuracy, stability, and switching costs in shaping policy performance. Conducted in the context of battery scheduling within energy management applications, it introduces a metric for evaluating probabilistic forecast stability and examines the effects of forecast accuracy and stability on optimization outcomes using the real-world case of the Citylearn 2022 competition. Findings indicate that switching costs significantly influence the trade-off between forecast accuracy and stability, highlighting the importance of integrated systems that enable collaboration between forecasting and operational units for improved decision-making. The study shows that committing to a policy for longer periods can be advantageous over frequent updates. Results also show a correlation between forecast stability and policy performance, suggesting that stable forecasts can mitigate switching costs. The proposed framework provides valuable insights for energy sector decision-makers and forecast practitioners when designing the operation of an energy management system.
@@
{highlights}
The paper evaluates the impact of forecast stability and accuracy on the downstream performance of energy management systems under switching costs.
A novel metric, Scenario Distribution Change (SDC), is proposed to assess the stability of probabilistic scenario sets, extending the Wasserstein distance.
Forecast accuracy and stability are shown to influence the optimal policy commitment level. There is a theoretical trade-off between switching costs and forecast errors.
Empirical analysis is conducted using the CityLearn 2022 competition case study, highlighting the trade-offs between forecast updates, stability, and operational outcomes. The longer than minimal commitment periods are found to improve system performance in case of stochastic optimization with probabilistic scenario sets.
@@ 1 Introduction
To effectively navigate these challenges, it is beneficial to integrate forecasting of the problem environment with the optimization of decision-making processes. This integration, often referred to as the predict, then optimize approach, is prevalent in energy applications like electric vehicle charging [1, 2], battery scheduling [3, 4], and energy and flexibility dispatch problems [5, 6]. In many of these applications, decision-makers face switching costs — expenses incurred from updating operational plans. For example, transitioning operations from one state to another often involves costs due to physical limitations on ramping energy production up or down [7], managing server load balancing [8], or covering network fees associated with energy trading [9, 10].
Despite extensive research on online optimization with predictions, there remains a notable gap concerning how switching costs affect the synergy between forecasting models and optimization algorithms. This study addresses this gap by exploring the concepts of time-coupling and forecast stability within such systems. Time-coupling, where one decision directly affects subsequent decisions and system states, gives rise to switching costs. Forecast inaccuracies often lead to policy revisions, requiring adjustments to better align with the current system state and updated information. Striking a balance between forecast accuracy and the impact of switching costs is crucial. This paper explores how this balance shapes system design and informs operational strategies. Consequently, this research aims to answer the following questions:
  • 1.
    Do switching costs influence the balance between forecast accuracy and stability?
  • 2.
    What are the optimal strategies for forecast revision in the presence of such costs?
  • 3.
    How can integrated systems be designed to effectively balance these considerations for improved decision-making in energy management?
@@ 1.1 Related work
Literature solutions for integrating forecasting and optimization are classified based on the linkage of these subproblems. There are three categories: Direct, Indirect, and Semi-direct methods.
@@
Direct methods engage with the optimization problem during training of the forecast making it an integrated approach which is known as ‘predict and optimize’. Such approach is proposed in Elmachtoub and Grigas [11], with optimization cost integrated directly into the forecast loss function to align forecasts with end-use optimization. This optimizes the prediction model in the forecasting stage for better decision-making in the optimization stage. The computational burden is partially resolved by using a simplified loss function, however more complex problems can still pose significant computational challenges. There are also other works that design an alternative task-specific loss function, such as reported in Mandi et al. [12], and tested on a real-world energy management problem.
@@
Indirect methods regard the two subproblems as separate, also referred in literature as ‘predict, then optimize’. This approach is regarded as standard and is widely applied in practice. In the study by Vanderschueren et al. [13], an empirical evaluation of direct and indirect approaches is conducted in the context of cost-sensitive classification. Although the study is not directly applicable to the scheduling problems, it is interesting to note that those authors find that the indirect approach where optimization is performed on the predictions of the classifier outperforms the integrated approach. Therefore, the effectiveness of cost-sensitive classification may not always align with the intuitively expected benefits of integrated training within the ’predict and optimize’.
@@
The third category, Semi-direct methods, considers characteristics of the optimization problem but abstains from direct interaction during training of the forecast. These methods can take various forms, as potential improvements can occur at any stage outside of the training process. An example is found in the work by Kazmi and Paskevich [14], where those authors propose a Bayesian Optimization approach for integrating the downstream optimization problem into the hyperparameter tuning process of the forecast model. Other example is [3], where the authors propose iterating attention over training data statically and dynamically to improve the forecast model performance in the downstream task. Among the semi-direct methods, there is a noticeable gap in the literature regarding the operational aspects of integrating forecasting models and optimization deployment. A relevant study by Prat et al. [15] addresses the issue of determining the minimum forecast horizon for storage scheduling problems in a rolling-horizon approach. The study introduces a verifiable condition to check if a chosen planning horizon is sufficiently long. Other researchers study the effect of the frequency of forecast revision, which can be interpreted as both the frequency of the forecast model deployment and the frequency of model training. The frequency of retraining is studied by Spiliotis and Petropoulos [16], where the authors investigate different scenarios of updating the model fit for univariate exponential smoothing (ES) and univariate gradient boosting models (LightGBM).
@@
We are particularly interested in studying the model deployment frequency. More specifically, we look at the applications of optimization with predictions in environments where switching costs are present. This topic has featured in research on online convex optimization (OCO) problems. In control theory, a widespread strategy for tackling online multi-step optimization challenges is the employment of the Receding Horizon Control (RHC) algorithm. Model Predictive Control (MPC) is the most common implementation of RHC that employs a dynamic model of the process and is widely used in online optimization. When rerun at every time step, this method commits to the first step of the future horizon while treating later decisions as advisory. The Fixed Horizon Control (FHC) algorithm is a generalization of the algorithm that commits to a fixed number of steps $v$ in the future before re-optimizing the policy. The difference between the two algorithms is illustrated in Figure 1. Throughout this paper, we will use the term FHC to align with the focus on commitment-specific optimization.
@@
The research conducted by Chen et al. [17] highlights the challenges posed by noisy predictions in designing Online Convex Optimization (OCO) algorithms. Several solutions, such as Averaging Fixed Horizon Control (AFHC) and its generalized counterpart, Committed Horizon Control (CHC) [18], propose aggregation strategies over multiple revisions of plans to achieve sublinear regret and improve robustness to prediction errors. Similarly, the Feasible Fixed Horizon Control (FFHC) algorithm [19] incorporates a regularization term to handle multi-interval optimization problems while ensuring feasibility and robustness against ramping costs.
While these approaches demonstrate the benefits of aggregating policies over prediction updates, the current work shifts focus to understanding the effects of commitment on forecast stability and downstream optimization in the presence of switching costs. Unlike these aggregation-based methods, this study seeks to explore how the commitment level impacts system performance, with a particular emphasis on the relationship between forecast stability and switching costs.
Forecast stability has been recognized as a critical factor in planning. Building on the comprehensive framework by Godahewa et al. [20], which categorizes stability into vertical (variations across forecast updates for the same period) and horizontal (variations within a prediction window), this work investigates these dimensions in the context of energy management systems.
Unlike supply chain management, where freezing intervals [21, 22] and smoothing techniques [23] have been widely used to address system nervousness and the bullwhip effect [24], energy systems require more dynamic and frequent forecast updates. Switching costs are also observed in energy management systems, such as ramping costs for generating units [7, 25] or toggling server load balancing [8, 26]. However, the literature lacks a systematic exploration of how commitment levels influence the interplay between forecast stability and switching costs. By focusing on these aspects, the current research sets the stage for advancing the design and operation of energy management systems, particularly in applications involving energy storage optimization and real-time decision-making.
@@ 1.2 Contents and Contributions
    First, we formally define the problem and the metric used for evaluating the downstream performance, both with and without switching costs. We also conduct a theoretical analysis of the performance bounds of the Fixed Horizon Control (FHC) algorithm, highlighting how the commitment level influences the trade-off between forecast accuracy and switching costs.
  • 2.
    The effect of forecast accuracy and stability on the downstream performance of the optimization policy is analyzed. We extend the definition of forecast stability towards probabilistic forecasts and propose a metric for measuring horizontal and vertical stability of scenario sets.
  • 3.
@@
    We conduct an empirical evaluation of the stability of point forecasts and probabilistic scenario sets, using deterministic and stochastic optimization approaches, respectively. We use the case study of the integrated forecasting and optimization energy management problem from the Citylearn 2022 competition. Based on the case study, we analyze the relationship between forecast accuracy, stability and the performance of the policy. We show that running the policy for a longer period provides a simple mechanism to avoid switching costs while maintaining or improving the performance.
  • 4.
    Lastly, the paper concludes with a discussion of the general implications of the results for the design of the energy management systems followed by suggested future research directions.
2 Methodology
2.1 Problem statement
Figure 2: Grid diagram illustrating the coordinated battery scheduling problem.
This paper addresses the challenge of optimizing battery storage management within a multi-agent energy management system. The setting of the problem is first outlined in the CityLearn Challenge 2022 [27] and based on a real-world dataset from a grid-interactive community. The CityLearn Challenge is an annual competition that focuses on the development of reinforcement learning agents for energy management in buildings. Participants are provided with a simulated environment that mimics real-world energy consumption and generation patterns. The goal is to optimize the operation of energy systems, such as batteries and HVAC systems, to minimize costs and emissions while maintaining comfort levels.
The series of CityLearn competitions, starting from 2020, was originally set up only for Reinforcement Learning (RL) solutions, where an agent learns to make decisions by performing actions and receiving feedback from the environment. However, from 2022 onwards, alternative methods such as Model Predictive Control (MPC), which uses a model to predict the future behavior of a system and optimize the control actions accordingly, and Rule-Based Control (RBC), which involves using predefined rules to make decisions based on the current state of the system, have also been employed. The competition provides a platform and the open-source simulation environment [^1] https://github.com/intelligent-environments-lab/CityLearn for implementing the operation of energy management systems with setups that include various energy assets, such as solar panels, batteries, heat pumps, connections to the grid, and electric vehicles.
The setup of the investigated problem is illustrated in Figure Figure 2. In general terms, we consider an online optimization problem with switching costs. The data of the competition has the hourly resolution. The core problem is multi-agent scheduling of energy storage over a period of a year. While the problem can be formulated similarly to RL, or a dynamic programming approach, in this paper, we particularly investigate the mechanics with a rolling-horizon approach. This means that the problem is solved iteratively over a planning horizon $H$. However, it is typical for only first $v$ steps to be implemented with the rest of the prediction horizon $H-v$ to be advisory, since the optimization is revised with a period of $v$ steps. The full details of the objective score and relevant constraints are provided in A.1. The problem includes finding an optimal state-dependent policy $X_{t}^{FHC}(S_{t}|\theta^{LA})$. The decision-maker is faced with a sequence of decisions, where the decision at each step is based on the current state $S_{t}$ of the system and the direct lookahead approximation $\theta^{LA}$. Using the framework proposed by Powell [28], we use the following formulation:
$X_{t}^{FHC}(S_{t}|\theta^{LA})=\arg\min_{x_{t}}\Biggl{(}\sum_{i\in n}W_{i}\cdot\mathbb{C}(\textbf{x};\textbf{S})\Biggr{)}=$
@@ 2.1 Problem statement
$+\beta\left\lVert\tilde{y}_{tt^{\prime}}-\tilde{y}_{t(t^{\prime}-1)}\right\rVert\Bigr{]}\Biggr{)}$
@@
  • 1.
    $X_{t}^{FHC}(S_{t}|\theta^{LA})$ Represents the optimal set of decisions (battery charging $x_{t}^{pos}$ and discharging $x_{t}^{neg}$ actions) made by the FHC at time $t$ based on the current state $S_{t}$ and the lookahead approximation model $\theta^{LA}$. This model enables predicting future costs within a horizon $H$, taking into account the expected electric load and PV generation.
  • 2.
    $\mathbb{C}(\textbf{x};\textbf{S})$ is the cost function, which is a function of actions $\textbf{x}=x_{t},...x_{t+H}$ and states $\textbf{S}=S_{t},...S_{t+H}$.
  • 3.
  • 4.
    $\mathit{h}(x_{tt^{\prime}};\theta^{LA}_{i,tt^{\prime}})$ Represents the cost of the set of decisions $x_{t+1},...,x_{t+H}$ given the future forecast scenarios generated by the model $\theta^{LA}_{i,tt^{\prime}}$
  • 5.
@@
Optimized battery schedules are evaluated using a set of Key Performance Indicators (KPIs). The optimization targets the minimization of the equally weighted sum of the normalized electricity cost $C$ and carbon emissions $G$. When grid-related metrics are included, the average includes the grid score $D$, which is computed as the mean of the normalized ramping KPI $R$ and the Load Factor KPI $(1-L)$. The grid score, $D$, introduces switching costs in the optimization by accounting for the ramping costs associated with frequent changes in forecasting. Therefore, in order to provide a comprehensive analysis of the impact of switching costs on the optimization, the combined optimization scores are defined with and without the grid score:
Average Score $=\operatorname{avg}\left(\frac{C_{\text{entry}}}{C_{\text{no\ battery}}},\frac{G_{\text{entry}}}{G_{\text{no\ battery}}}\right)$ (2)
Average Score (with switching costs) $=\operatorname{avg}\left(\frac{C_{\text{entry}}}{C_{\text{no\ battery}}},\frac{G_{\text{entry}}}{G_{\text{no\ battery}}},D\right)$ (3)
All the scores range from 0 to 1 because they are normalized by the no-battery scenario, representing the improvement as a fraction of the metric compared to the no-battery scenario. The full formulation of the cost function, along with the definition of each of the KPIs and the formulation of the MPC problem is given in A.2. Furthermore, In A.3, we conduct a theoretical analysis of the performance bounds on the solving algorithm in an online optimization problem with switching costs. Following the line of proof in Chen et al. [18], we derive the upper bound on the competitive difference between the FHC algorithm and the optimal policy. The analysis is conducted for the deterministic case, where the forecast is a point forecast, then extended to the probabilistic case, where the forecast is represented as a set of scenarios. We discuss the role of policy commitment and draw a parallel to stability. A trade off is demonstrated between forecast accuracy and switching costs, with the commitment level regulating the balance between the two.
2.2 Optimization
The problem is solved using the Fixed Horizon Control (FHC) algorithm, which is a configuration of model predictive control (MPC) with a fixed commitment horizon. The optimal policy is computed for the current state of the system and the forecast for the next $H$ time steps. An online optimization algorithm typically determines the optimal actions $x_{t+1},...,x_{t+H}$ given a prediction window of length $H$. Every $v$ time steps, the optimization is rerun to generate a new plan. This algorithm is referred to as a Fixed Horizon Control (FHC) algorithm.
In the context of the online optimization algorithms, the term “commitment” refers to the duration or number of time steps for which the optimized policy is committed to before the optimization process is rerun to generate a new plan. This concept plays a crucial role in determining how frequently the optimization is updated based on new information, accumulation of forecast errors and changes in the system’s state. An FHC algorithm with the shortest commitment is the Receding Horizon Control (RHC) algorithm. In RHC, the commitment is $v=1$, indicating that the optimization is rerun at every time step.
Similar to the optimization, the forecast is also run in a rolling origin fashion. The optimal policy is then applied to the system for the next $v_{F}$ time steps and the process is repeated every $v_{F}$ steps. In summary, both the forecast and the optimization have the same rolling origin setup and need to be updated at certain intervals while committing to $v$ steps. Ultimately, at each time step, the decision-maker is faced with a decision to either reuse the current plan or to revise it. Similarly, the forecast can be reused or revised. The resulting decision matrix is shown in Table 1.

### v5

# Balancing Forecast Accuracy and Switching Costs in Online Optimization of Energy Management Systems

@@ Abstract
This study investigates the integration of forecasting and optimization in energy management systems, with a focus on the role of switching costs, defined as penalties incurred from frequent operational adjustments. We develop a theoretical and empirical framework to examine how forecast accuracy and stability interact with switching costs in online decision-making settings. Our analysis spans both deterministic and stochastic optimization approaches, using point and probabilistic forecasts. A novel metric for measuring temporal consistency in probabilistic forecasts is introduced, and the framework is validated in a real-world battery scheduling case based on the CityLearn 2022 challenge. Results show that switching costs significantly alter the trade-off between forecast accuracy and stability, and that more stable forecasts can reduce the performance loss due to switching. Contrary to common practice, the findings suggest that, under non-negligible switching costs, longer commitment periods may lead to better overall outcomes. These insights have practical implications for the design of intelligent, forecast-aware energy management systems.
@@
[…]
[cor1]Corresponding author
@@ 1 Introduction
To effectively navigate these challenges, it is beneficial to integrate forecasting of the problem environment with the optimization of decision-making processes. This integration, often referred to as the predict, then optimize approach, is prevalent in energy applications like electric vehicle charging Kriekinge et al. [2021], Tungom et al. [2024], battery scheduling Ruddick et al. [2024], Liu et al. [2023], and energy and flexibility dispatch problems Qu et al. [2024], Liu et al. [2025]. In many of these applications, decision-makers face switching costs — expenses incurred from updating operational plans. For example, transitioning operations from one state to another often involves costs due to physical limitations on ramping energy production up or down Wu et al. [2013], managing server load balancing Gupta [2021], Lin et al. [2012], or covering network fees associated with energy trading Yang et al. [2023], Levorato et al. [2022].
Despite extensive research on online optimization with predictions, there remains a notable gap concerning how switching costs affect the synergy between forecasting models and optimization algorithms. This study addresses this gap by exploring the concepts of time-coupling and forecast stability within such systems. Time-coupling, where one decision directly affects subsequent decisions and system states, gives rise to switching costs. Forecast inaccuracies often lead to policy revisions, requiring adjustments to better align with the current system state and updated information.
A key consideration in these systems is determining the optimal commitment period, meaning how long one can commit to a given policy before recalculating based on updated forecasts. This commitment level directly influences system performance, especially when switching costs are present. Additionally, forecast stability, the consistency of predictions across successive updates, plays a critical role in the reliability and efficacy of the optimization process. Striking a balance between forecast accuracy, stability, and the impact of switching costs is crucial. This paper explores how this balance shapes system design and informs operational strategies. Consequently, this research aims to answer the following questions:
  • Does presence of switching costs affect the optimal commitment level of the policy?
  • How do forecast accuracy and stability influence the downstream performance of energy management systems?
  • How can integrated systems of forecasting and optimization be designed to effectively balance these considerations for improved decision-making in energy management?
@@ 1.1 Related work
Literature solutions for integrating forecasting and optimization are classified based on the level of integration between the two components. This classification includes three categories: Direct, Indirect, and Semi-direct methods. Direct methods engage with the optimization problem during training of the forecast making it an integrated approach which is known as ‘predict and optimize’. Such approach is proposed in Elmachtoub and Grigas [2022], with optimization cost integrated directly into the forecast loss function to align forecasts with end-use optimization. This optimizes the prediction model in the forecasting stage for better decision-making in the optimization stage. The computational burden is partially resolved by using a simplified loss function, however more complex problems can still pose significant computational challenges. There are other works that design an alternative task-specific loss function, such as reported in Mandi et al. [2020], and tested on a real-world energy management problem.
@@
Indirect methods regard the two sub-problems as separate, also referred in literature as ‘predict, then optimize’. This approach is regarded as standard and is widely applied in practice. In the study by Vanderschueren et al. [2022], an empirical evaluation of direct and indirect approaches is conducted in the context of cost-sensitive classification. Although the study is not directly applicable to the scheduling problems, it is interesting to note that those authors find that the indirect approach where optimization is performed on the predictions of a classification algorithm outperforms the integrated approach. Therefore, the effectiveness of cost-sensitive classification may not always align with the intuitively expected benefits of integrated training within the ’predict and optimize’.
@@
The third category, Semi-direct methods, considers characteristics of the optimization problem but abstains from direct interaction during training of the forecast. These methods can take various forms, as potential improvements can occur at any stage outside the training process. An example is found in the work by Kazmi and Paskevich [2023], where those authors propose a Bayesian Optimization approach for integrating the downstream optimization problem into the hyperparameter tuning process of the forecast model. Other example is Huang et al. [2025], where the authors propose iterating attention over training data statically and dynamically to improve the forecast model performance in the downstream task. Among the semi-direct methods, there is a noticeable gap in the literature regarding the operational aspects of integrating forecasting models and optimization deployment. A relevant study by Prat et al. [2024] addresses the issue of determining the minimum forecast horizon for storage scheduling problems in a rolling-horizon approach. The study introduces a verifiable condition to check if a selected planning horizon is sufficiently long. Other researchers study the effect of the frequency of forecast revision, which can be interpreted as both the frequency of the forecast model deployment and the frequency of model training. The frequency of retraining is studied by Spiliotis and Petropoulos [2024], where the authors investigate different scenarios of updating the model fit for univariate exponential smoothing (ES) and univariate gradient boosting models (LightGBM Ke et al. [2017]).
@@
We are particularly interested in studying the model deployment frequency. More specifically, we look at the applications of optimization with predictions in environments where switching costs are present. This topic has featured in research on online convex optimization (OCO) problems. In control theory, a widespread strategy for tackling online multi-step optimization challenges is the employment of the Receding Horizon Control (RHC) algorithm. When rerun at every time step, this method commits to the first step of the future horizon while treating later decisions as advisory. The Fixed Horizon Control (FHC) algorithm is a generalization of the algorithm that commits to a fixed number of steps $v$ in the future before re-optimizing the policy. The difference between the two algorithms is illustrated in Figure 1. Throughout this paper, we will use the term FHC to align with the focus on commitment-specific optimization. In literature, the FHC/RHC algorithm is often referred to as Model Predictive Control (MPC), which is a well-established method for solving online multi-step optimization problems in control theory.
@@
The challenge of noisy predictions in Online Convex Optimization (OCO) has been extensively studied, notably by Chen et al. [2015], who emphasized the difficulty of designing robust online algorithms under uncertainty. In response, various strategies have emerged to mitigate the impact of prediction noise. For instance, Averaging Fixed Horizon Control (AFHC) and its generalized form, Committed Horizon Control (CHC) Chen et al. [2016], employ aggregation across multiple plan revisions to achieve sublinear regret and greater robustness. Likewise, the Feasible Fixed Horizon Control (FFHC) algorithm Christianson et al. [2022] introduces regularization to ensure feasibility in multi-interval settings while addressing ramping-related switching costs. In this study, we adopt the classical Fixed Horizon Control (FHC) algorithm with a fixed commitment level as the foundational optimization scheme. FHC is not only conceptually simple but also widely adopted in energy management systems, particularly in implementations of Model Predictive Control (MPC) Ahmad et al. [2023]. While previous works have primarily focused on aggregating or averaging policies to improve performance, our focus shifts toward examining how the frequency of re-optimization, i.e., the commitment level, interacts with forecast stability and switching costs.
Forecast stability itself has recently gained attention as a key dimension in sequential decision-making. Building on the taxonomy proposed by Godahewa et al. [2023], which distinguishes between vertical stability (variation across forecast updates for the same time step) and horizontal stability (variation within the forecast horizon at a single time step), we explore how these notions influence system performance in the energy domain. Though similar concerns have been addressed in supply chain literature, where freezing intervals Sahin et al. [2013], Xie et al. [2003] and smoothing techniques Ho* [2005] aim to reduce system nervousness and the bullwhip effect Wang and Disney [2016], energy systems impose different requirements. Forecasts must be updated more frequently and switching costs (such as ramping costs Wu et al. [2013], Tanaka [2006] or toggling server states Lin et al. [2012], Lee et al. [2021]) are often tied to physical infrastructure.
Despite the relevance of these factors, few studies have systematically explored the link between commitment strategies, forecast stability, and switching costs in energy systems. This study addresses that gap by investigating how varying commitment levels affect performance, particularly in energy storage scheduling and real-time optimization contexts. By doing so, we aim to inform the design of more robust and responsive decision-making frameworks in energy management.
@@ 1.2 Contents and Contributions
  • First, we formally define the problem and the framework for online optimization with switching costs, introducing the Fixed Horizon Control (FHC) algorithm and the concept of forecast stability. We develop the theoretical analysis that highlights how the commitment level influences the trade-off between forecast accuracy and switching costs.
  • We propose a novel metric called Scenario Distribution Change (SDC) for evaluating the stability of probabilistic scenario sets, extending the concept of forecast stability from point forecasts to probabilistic forecasts. This metric allows for measuring both horizontal stability (within a forecast horizon) and vertical stability (across forecast updates).
@@
  • We conduct an empirical evaluation using the integrated forecasting and optimization energy management problem from the Citylearn 2022 competition. Through extensive testing of different commitment periods, we analyze the relationship between forecast accuracy, stability, and policy performance, demonstrating that in the presence of switching costs, longer periods of commitment can lead to improved decision-making and system performance.
  • Lastly, we discuss the implications of our findings for the design of energy management systems, highlighting how the balance between forecast accuracy, stability, and switching costs shapes optimal operational strategies in real-world applications.
2 Problem Formulation and Framework
This section begins by introducing a general problem framework for online optimization with switching costs, defining the mathematical formulation and related key concepts. Next, we present the Fixed Horizon Control algorithm as the primary solution method, discussing how commitment periods affect decision-making. The concept of forecast stability is also introduced to analyze its impact on system performance.
2.1 Online Optimization with Switching Costs
Online optimization with switching costs is a class of problems characterized by sequential decision-making under uncertainty. In these problems, decisions must be made in real-time as new information becomes available, with only partial knowledge of future conditions. What makes this class of problems particularly challenging is the presence of time-coupling effects, where decisions at one time step directly affect the state of the system and the options available at subsequent time steps. The general problem is defined as finding an optimal state-dependent policy $X_{t}(S_{t}|\theta^{LA})$. The decision-maker faces a sequence of decisions, where each decision is based on the current system state $S_{t}$ and the direct look-ahead approximation $\theta^{LA}$. Using the framework proposed by Powell [2019], the problem can be formulated as follows:
$X_{t}(S_{t}|\theta^{LA})=\arg\min_{x_{t}}\Biggl{(}\sum_{i\in n}W_{i}\cdot\mathbb{C}(\textbf{x};\textbf{S})\Biggr{)}=$
@@ 2.1 Online Optimization with Switching Costs
$+\beta\left\lVert\tilde{S}_{tt^{\prime}}-\tilde{S}_{t(t^{\prime}-1)}\right\rVert\Bigr{]}\Biggr{)}$
@@
  • $\mathbb{C}(\textbf{x};\textbf{S})$ denotes the total cost function, which is a function of the decision variables x and the system state S
  • $W_{i}$ represents the weight associated with different cost components or scenarios. For deterministic optimization, only one scenario exists, while stochastic approaches consider multiple weighted scenarios.
  • $\mathit{h}(x_{tt^{\prime}};\theta^{LA}_{i,tt^{\prime}})$ captures the direct operational cost at time $t^{\prime}$ when following decision $x_{tt^{\prime}}$ under forecast scenario $\theta^{LA}_{i,tt^{\prime}}$.
@@
2.2 Fixed Horizon Control Algorithm
To address online optimization problems with switching costs, the Fixed Horizon Control (FHC) algorithm is employed. FHC is a variant of model predictive control with a fixed commitment horizon. At each decision point, the algorithm computes an optimal policy for the current state of the system and forecasts for the next $H$ time steps, determining the optimal actions $x_{t+1},...,x_{t+H}$ given a prediction window of length $H$.
A key parameter in FHC is the commitment period $v$, which defines how many steps of the computed policy are implemented before recomputing. Every $v$ time steps, the optimization is rerun to generate a new plan. When $v=1$, the algorithm becomes Receding Horizon Control (RHC), where optimization is performed at every time step, implementing only the first step of each plan. Both the forecast and optimization follow a rolling origin setup and require updates at certain intervals while committing to $v$ steps. At each time step, the decision-maker faces a choice to either reuse the current plan or to revise it. Similarly, the forecast can be reused or revised. This decision framework is illustrated in Table 1. It is important to note that not all quadrants in this decision matrix are equally practical or beneficial. Specifically, updating the forecast without updating the optimization plan (bottom-left quadrant) offers limited value, as new information is obtained but not acted upon. Conversely, updating the optimization without updating the forecast (top-right quadrant) can be valuable in situations where the system state changes in ways unrelated to the forecast variables. In most practical implementations, the most logical configurations are either reusing both forecast and plan (top-left) or revising both simultaneously (bottom-right).